OpenStudy (anonymous):

Help please: If a and b are +ve real numbers such that a+b=1,prove A^ab^b + a^bb^a ≤ 1

OpenStudy (anonymous):

Do you mean this for the second equation? \[a^a*b^b+a^b*b^a \le1\]

OpenStudy (anonymous):

yes, u r right...thanks

OpenStudy (anonymous):

b=1-a so f(a)=a^a(1-a)^1-a+a^(1-a) (1-a)^a

OpenStudy (anonymous):

are you sure it is right?

OpenStudy (anonymous):

mahmit yes...thanks

OpenStudy (anonymous):

it is an olympiad question

OpenStudy (anonymous):

find f(0+)=f(1-) then show it is 1 and both are corner maximum value.

OpenStudy (anonymous):



OpenStudy (anonymous):